OpenStudy (anonymous):

What is the 3rd term of the sequence?

OpenStudy (anonymous):

\[a_{1}=40 \] and \[a _{n}=0.5(a_{n}-1)\]

OpenStudy (anonymous):

\[a_{3}=\]

ganeshie8 (ganeshie8):

\[\rm a_1 = 40\] and \[a_n = 0.5a_{n-1}\] right ?

ganeshie8 (ganeshie8):

which is same as : \[a_n = \frac{1}{2}a_{n-1}\]

ganeshie8 (ganeshie8):

plugin n=2 and find the second term first

ganeshie8 (ganeshie8):

which is same as : \[a_2 = \frac{1}{2}a_{2-1} = \frac{1}{2} a_1 = \frac{1}{2}(40) = 20\]

ganeshie8 (ganeshie8):

use this to find \(\rm a_3\)

OpenStudy (anonymous):

400?

ganeshie8 (ganeshie8):

It seems you're taking half of present term to get the next term right ?

ganeshie8 (ganeshie8):

\[a_3 = \frac{1}{2}a_{3-1} = \frac{1}{2} a_2 = ?\]

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

@ganeshie8